# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 1.5 Approximate projections
  • iii)
    If a sequence $\{z^{k}\}\subset\mathds{R}^{n}$ converges to $z^{*}\in\mathds{R}^{n}$ and $\lim_{k\to\infty}{\rm dist}(z^{k},S(z^{k}))=0$, then $z^{*}\in K$,
  • iv)
@@
    ${\rm dist}\left(P_{S(x)\cap V}(x),K\right)\leq\tau\,{\rm dist}(x,K),$ (1.17)
where dist(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the Euclidean distance between a point and a set in $\mathds{R}^{n}$. This notion of separating set was introduced in [3], where only items (i)-(iii) in Definition 1.2 were demanded. In this paper, we additionally require item (iv). We commit a slight abuse of notation, keeping the same definition after adding (iv), which prevents the separating set $S(x)$ from being too close to $x$.
As long as we determine a separating set $S_{i}$ of $C_{i}$ that satisfies the assumptions in Definition 1.2, the exact projection onto $S_{i}$ can be regarded as the approximate projection onto the original set $C_{i}$. Suppose that $S_{i}(x)\subset\mathds{R}^{n}$ is a separating set of $C_{i}$ at point $x\in\mathds{R}^{n}$ for each $i\in\{1,\ldots,m\}$. In order to utilize CRM with approximate projection, we first define the separating set in the product space as follows:
$\mathbf{S}:=S_{1}\times\cdots\times S_{m}\subset\mathds{R}^{nm}.$ (1.18)
@@
$\mathbf{T}_{\mathbf{S},\mathbf{D}}(\mathbf{z})={\rm circ}(\mathbf{R}_{\mathbf{D}}\mathbf{R}_{\mathbf{S}}(\mathbf{z}),\mathbf{R}_{\mathbf{S}}(\mathbf{z}),\mathbf{z}),$ (1.19)
where $circ(\cdot,\cdot,\cdot)$ are the circumcenter of their points as it defined in (1.11), $\mathbf{R}_{\mathbf{S}}$ is defined in the same as in (1.15) but with respect to $\mathbf{S}$ defined in (1.18) and $\mathbf{R}_{\mathbf{D}}$ is defined in (1.16). Therefore, the sequence $\{\mathbf{z}^{k}\}\subset\mathds{R}^{nm}$ generated by ACRM can be written as
$\mathbf{z}^{k+1}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{z}^{k}),$ (1.20)
@@
and $S_{i}^{k}\subset\mathds{R}^{n}$ is any separating set of $C_{i}$ at point $x^{k}$, and $\mathbf{z}^{k}=(x^{k},\ldots,x^{k})$. In view of Proposition 2.7(i) in [5], we conclude that $\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}:\mathbf{D}:\rightarrow\mathbf{D}$, that is $\{\mathbf{z}^{k}\}\subset\mathbf{D}$ for all $k\geq 1$. Now we go back to $\mathbf{\mathds{R}}^{n}$, and define the approximate circumcentered-reflection operator $T_{\mathbf{S}_{k},\mathbf{D}}:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ as follows
$T_{\textbf{S}_{k},\textbf{D}}(x)=\bar{x},$ (1.21)
@@
$\mathbf{\bar{x}}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}),\qquad\mathbf{\bar{x}}=(\bar{x},\dots,\bar{x})\qquad\mathbf{x}=(x,\dots,x),$ (1.22)
@@
$x^{k+1}=T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}).$ (1.23)
In practice, the sets $S^{k}_{i}$ will be halfspaces, so that the projections onto the sets $S^{k}_{i}$ are easily computable. The same holds for the operators $R_{\textbf{W}},R_{{\textbf{S}}_{k}}$, and consequently for the operator $T_{{\textbf{S}}_{k},{\textbf{D}}}$. This allows us to obtain a closed formula for $x^{k+1}$ in terms of $x^{k}$ in (1.21). This is the case, for instance, of Algorithm PACA (see [11]), which, using suitably perturbed separating sets, generates a sequence that is finitely convergent to a solution of CFP, under a Slater condition.
2 Preliminaries
@@
Assume that $K\subset\mathds{R}^{n}$ is a closed convex set. Then,
  • i)
    $\langle x-P_{K}(x),y-P_{K}(x)\rangle\leq 0,~{}\forall x\in\mathds{R}^{n}~{}and~{}\forall y\in K.$
  • ii)
@@
If $H=\{y\in\mathds{R}^{n}\mid a^{\top}y\leq\alpha\}$, with $a\in\mathds{R}^{n}$, $\alpha\in R$, then the orthogonal projection onto $H$ is given by
$P_{H}(x)=x-\frac{\max\{0,\alpha-a^{\top}x\}}{\|a\|^{2}}a.$
@@
Suppose $C\subset\mathds{R}^{n}$ is a closed convex set, $U\subset\mathds{R}^{n}$ is an affine subspace. Let $S\subset\mathds{R}^{n}$ be a separating set of $C$ which satisfies $(i)-(iii)$ in Definition 1.2. Define $H(x)\subset\mathds{R}^{n}$ as:
$H(x)=\begin{cases}K,&{\rm if}\,\,\,x\in C,\\ \{z\in\mathds{R}^{n}:\langle z-P_{S}(x),x-P_{S}(x)\rangle\leq 0\},&{\rm otherwise}.\end{cases}$ (2.1)
@@
Next, we present an explicit instance of the approximate circumcenter operator $T_{\mathbf{S},\mathds{D}}$, defined in (1.21). First, we define a separating set $S_{i}(y)\subset\mathds{R}^{n}$ associated to each set $C_{i}$, assumed to be of the form:
$C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$ (2.3)
@@
$S_{i}(y)=\begin{cases}C_{i}&{\rm if}\,\,y\in C_{i},\\ \{z\in\mathds{R}^{n}:(u_{i})^{\top}(z-y)+g_{i}(x)\leq 0\},&{\rm otherwise},\end{cases}$ (2.4)
where $u_{i}$ is a subgradient of $g_{i}$ at $x^{k}$. Given ${\mathbf{y}}=(y^{1},\dots,y^{m})\in\mathds{R}^{nm}$, we define ${\mathbf{S}}({\mathbf{y}})\subset\mathds{R}^{nm}$ as
${\mathbf{S}}({\mathbf{y}})=S_{1}(y^{1})\times\dots\times S_{m}(y^{m}).$ (2.5)
We have the following result related to these separating sets.
@@
Suppose that $C_{i}\subset\mathds{R}^{n}$ and $S_{i}(y)\subset\mathds{R}^{n}$ are defined as in (2.3) and (2.4). Then $C_{i}\subset S_{i}(y)$ for any $y\in\mathds{R}^{n}$ and $S_{i}(y)$ is a separating set for $C_{i}$ at the point $y$ that satisfies conditions (i)-(iii) in Definition 1.2. Also, ${\mathbf{S}}({\mathbf{y}})$ is a separating set for $C_{1}\times\dots\times C_{m}\subset\mathds{R}^{nm}$ at ${\mathbf{y}}$.
@@
Take $C_{i}$ as in (2.3), $T_{\mathbf{S}_{k},\mathbf{D}}$ as in (1.21) and $\mathbf{S}_{k}$ as above. By Proposition 2.4 and 2.3, we immediately have $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)=P_{\mathbf{H}_{k}\cap\mathbf{D}}(x)$, where $\mathbf{H}_{k}$ is defined in the same way as $H(x)$ in (2.1) with respect to the separating set $\mathbf{S}_{k}$.
@@
$\left\lVert T_{K,D}(x)-y\right\rVert\leq\left\lVert T_{MAP}(x)-y\right\rVert,$ (2.6)
for all $x\in U$, for all $y\in K\cap U$ and $T_{K,D}$ is the circumcenter operator as defined in (1.11) and $T_{MAP}=P_{U}\circ P_{K}$ is the alternating projection operator with respect to $K$ and $U$.
@@
If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is continuous then SOL$(T,C)$ is closed.
@@
Follows the definition of VIP$(F,C)$ and the continuity of $F$. ∎
@@
Let $F:\mathds{R}^{n}\to\mathds{R}^{n}$ be a continuous and paramonotone operator. Let $\{z^{k}\}$ be a bounded sequence such that all cluster points of $\{z^{k}\}$ belong to $C$. For each $x\in{\rm SOL}(F,C)$ define $\gamma_{k}:=\langle F(x^{k}),z^{k}-x\rangle$. If for some $\bar{x}\in{\rm SOL}(F,C)$ there exists a subsequence $\{\gamma_{k_{j}}\}$ of $\{\gamma_{k}\}$ such that $\lim_{j\to\infty}\gamma_{k_{j}}\leq 0$, then there exists a cluster point of $\{z^{k_{j}}\}$ belonging to ${\rm SOL}(F,C)$.
@@
Take $\{\varepsilon_{k}\}$, $\{v_{k}\}$subset $\mathds{R}_{+}$ and $\mu\in[0,1)$. If the inequalities
$\varepsilon_{k+1}\leq\mu\varepsilon_{k}+v_{k},~{}~{}k\in\mathds{N}$ (2.7)
hold and $\lim_{k\rightarrow\infty}v_{k}=0$, then $\lim_{k\rightarrow}\varepsilon_{k}=0$.
@@
${\rm dist}(z^{k_{j}+1},K)>{\rm dist}(z^{k_{j}},K)$ (2.8)
@@
We say $w\in\mathds{R}^{n}$ is a Slater point for ${\rm SOL}(F,C)$ if
$g_{i}(w)<0,$ (2.10)
@@
Let $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ be monotone and continuous and $C$ a closed and convex set. Then ${\rm SOL}(F,C)$ is closed and convex.
@@
${\rm SOL}(F,C)=\{x\in C:\langle F(y),y-x\rangle\geq 0,~{}\forall y\in C\}.$ (2.12)
@@
If $\{x^{k}\}$ is a quasi-Fejér sequence with respect to $S$ then:
  • (i)
@@
  • (i)
    the sequence $\{dist(x^{k},S)\}$ is convergent;
  • (ii)
@@
Let $g_{i}:\mathds{R}^{n}\rightarrow\mathds{R}$ is convex function for all $i\in\{1,\ldots,m\}$ and $C_{i}:=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Define $C:=\cap_{i=1}^{m}C_{i}$. Assume that there exists a Slater point $w\in C$ such that $g(w)<0$ where $g$ is defined in (4.5). Then, for all $y$ such that $g(y)>0$, we have
$\operatorname*{dist}(y,C)\leq\frac{\|y-w\|}{g(y-g(w))}g(y).$ (2.14)
@@
Take $w_{\lambda}:=\lambda w+(1-\lambda)y$ with $\lambda=\frac{g(y)}{g(y)-g(w)}$. Note that $\lambda\in(0,1)$. Then
$g(w_{\lambda})=g(\lambda w+(1-\lambda)y)\leq\lambda g(w)+(1-\lambda)g(y)=g(y)-\lambda(g(y)-g(w))=0,$ (2.15)
@@
$dist(y,C)\leq\|y-w_{\lambda}\|=\|y-(\lambda w+(1-\lambda)y)\|=\lambda\|y-w\|=\frac{g(y)}{g(y)-g(w)}\|y-w\|.$ (2.16)
3 A circumcenter method for the variational inequality problem with paramonotone operator
In this paper, we focus on the variational inequality problem VIP$(F,C)$ with a feasible set of the form $C:=\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex, i.e., we deal with the problem:
${\rm find}~{}x^{*}\in C\\ \,\,\,{\rm such\,\,\,that}\,\,\,\langle F(x^{*}),x-x^{*}\rangle\geq 0\,\,\,\forall x\in C.$ (3.1)
We propose two different algorithms, depending on the monotonicity and continuity properties of $F$. Therefore, we list next several assumptions, some of which will be needed for the convergence analysis of each algorithm.
Assumption 3.1.
The solution set of Problem (3.1), i.e., SOL$(F,C)$, is nonempty.
Assumption 3.2.
The mapping $F$ is Lipschitz continuous on $C$ with some constant $L>0$, i.e.,
$\left\lVert F(x)-F(y)\right\rVert\leq L\left\lVert x-y\right\rVert,\,\,\,\forall x,y\in C.$ (3.2)
Assumption 3.3.
The mapping $F$ is monotone on $C$.
Assumption 3.4.
The mapping $F$ is paramonotone on $C$.
Assumption 3.5.
Some comments are in order regarding Assumption 3.5. Coerciveness assumptions have been required in previous convergence analysis of direct methods for variational inequalities, for instance in [33], where a condition stronger than Assumption 3.5 is demanded, and in [17], where Assumption 3.5 is used. Our assumption holds immediately when $F$ is monotone and has a zero $\bar{z}$, in which case we take $\hat{z}=\bar{z}$ and $D=\emptyset$, and also when $F$ is uniformly monotone with $\lim_{t\to\infty}\psi(t)/t=\infty$ (for more details, see [17], where other monotone operators satisfying this assumption are discussed).
In this section, we consider a one-step circumcenter method, based on the GT method, for solving VI$(F,C)$ with a paramonotone operator $F$ and a feasible set of the form $C=\cap_{i=1}^{m}$.
The starting point for our method will be Algorithms 1 and 2 in [17], which are versions of the TG method for point-to-set paramonotone operators. We will add the circumcenter approach introduced in [14] and [15].
3.1 The algorithm
We present next the one-step circumcenter method for solving VIP$(F,C)$, with paramonotone $F$ and feasible set $C$ of the form $C=\cap_{i=1}^{m}C_{i}$.
Algorithm 1 Approximate Circumcenter method for VIP$(F,C)$ with paramonotone $F$
Initialization step
: Choose $x^{0}\in\mathds{R}^{n}$.
Iterative step
: Given $x^{k}$, if $F(x^{k})=0$ then stop. Otherwise, define $\eta_{k}:=\max\{1,\left\lVert F(x^{k})\right\rVert\}$. Choose a separating operator $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ for the set $C_{i}$, define $S^{i}_{k}:=S^{i}(x^{k})$ and ${\mathbf{S}}_{k}:=S^{1}_{k}\times\dots\times S^{m}$.Then, define the next iterate as:
$x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right),$ (3.3)
where $T_{\textbf{S}_{k},\textbf{D}}$ is the approximate circumcentered-reflection operator defined in (1.23), and the stepsize sequence $\{\beta_{k}\}$ satisfies that:
$\sum_{k=0}^{\infty}\beta_{k}=\infty,$ (3.4)
and
$\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty,$ (3.5)

### v2

@@ 1.5. Approximate projections
  3. (iii)
    If a sequence $\{z^{k}\}\subset\mathds{R}^{n}$ converges to $z^{*}\in\mathds{R}^{n}$ and $\lim_{k\to\infty}\operatorname*{dist}(z^{k},S(z^{k}))=0$, then $z^{*}\in K$,
  4. (iv)
@@
    (1.14) $\operatorname*{dist}\left(P_{S(x)\cap V}(x),K\right)\leq\tau\,\operatorname*{dist}(x,K),$
where $\operatorname*{dist}(\cdot,\cdot)$ denotes the Euclidean distance between a point and a set in $\mathds{R}^{n}$.
This notion of separating operator was introduced in [3], where only items (i)-(iii) in Definition 1.4 were demanded. In this paper, we additionally require item (iv). We commit a slight abuse of notation, keeping the same definition after adding (iv), which prevents $S(x)$ from being too close to $K$. As long as we determine a separating operator $S_{i}$ for $C_{i}$ that satisfies the assumptions in Definition 1.4, the exact projection onto $S_{i}$ can be regarded as the approximate projection onto the original set $C_{i}$. Suppose that $S_{i}(x)\subset\mathds{R}^{n}$ is a separating operator for $C_{i}$ at point $x\in\mathds{R}^{n}$ for each $i\in\{1,\ldots,m\}$. In order to utilize CRM with approximate projection, we first define the separating operator in the product space as follows:
(1.15) $\mathbf{S}\coloneqq S_{1}\times\cdots\times S_{m}\subset\mathds{R}^{nm}.$
@@
(1.16) $\mathbf{T}_{\mathbf{S},\mathbf{D}}(\mathbf{z})=\operatorname*{circ}(\mathbf{R}_{\mathbf{D}}(\mathbf{R}_{\mathbf{S}}(\mathbf{z})),\mathbf{R}_{\mathbf{S}}(\mathbf{z}),\mathbf{z}),$
where $\operatorname*{circ}(\cdot,\cdot,\cdot)$ is the circumcenter operator defined in (1.8), $\mathbf{R}_{\mathbf{S}}$ is defined in the same way as in (1.12) but with respect to $\mathbf{S}$ defined in (1.15) and $\mathbf{R}_{\mathbf{D}}$ is defined in (1.13). Therefore, the sequence $\{\mathbf{z}^{k}\}\subset\mathds{R}^{nm}$ generated by ACRM can be written as
(1.17) $\mathbf{z}^{k+1}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{z}^{k}),$
@@
$S_{i}^{k}$ is short for $S_{i}(x^{k})$ which is any separating operator for $C_{i}$, and $\mathbf{z}^{k}=(x^{k},\ldots,x^{k})$. In view of Proposition 2.7(i) in [4], we conclude that $\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}:\mathbf{D}:\rightarrow\mathbf{D}$, that is $\{\mathbf{z}^{k}\}\subset\mathbf{D}$ for all $k\geq 1$. Now we go back to $\mathbf{\mathds{R}}^{n}$, and define the approximate circumcentered-reflection operator $T_{\mathbf{S}_{k},\mathbf{D}}:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ as follows
(1.18) $T_{\textbf{S}_{k},\textbf{D}}(x)=\bar{x},$
@@
(1.19) $\mathbf{\bar{x}}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}),\qquad\mathbf{\bar{x}}=(\bar{x},\dots,\bar{x})\qquad\mathbf{x}=(x,\dots,x),$
@@
(1.20) $x^{k+1}=T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}).$
In practice, the sets $S^{k}_{i}$ will be halfspaces, so that the projections onto the sets $S^{k}_{i}$ are easily computable. The same holds for the operators $R_{\textbf{W}},R_{{\textbf{S}}_{k}}$, and consequently for the operator $T_{{\textbf{S}}_{k},{\textbf{D}}}$. This allows us to obtain a closed formula for $x^{k+1}$ in terms of $x^{k}$ in (1.18). This is the case, for instance, of Algorithm PACA (see [10]), which, using suitably perturbed separating operators, generates a sequence that is finitely convergent to a solution of CFP, under a Slater condition.
In this paper, we focus on the variational inequality problem $\operatorname{VIP}(F,C)$ with a feasible set of the form $C\coloneqq\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex, i.e., we deal with the problem:
(1.21) ${\rm find}~x^{*}\in C\\ \,\,\,{\rm such\,\,\,that}\,\,\,\langle F(x^{*}),x-x^{*}\rangle\geq 0\,\,\,\forall x\in C.$
1.6. Our contributions and assumptions
In this work we propose two different algorithms, depending on the monotonicity and continuity properties of $F$, that make use of the circumcenter machinery to accelerate convergence. The firt one (Algorithm 1) is an acceleration of the algorithm proposed in [17], while the second one (Algorithm 2) accelerates an algorithm proposed in [18] for monotone operators. Both algorithms (and our novel scheme as well) use approximate projections onto separating halfspaces of the sets $C_{i}$ instead of exact projections onto $C$. We prove that our methods generate sequences converging to a solution of $\operatorname{VIP}(F,C)$ under mild assumptions.
We next list several assumptions, which will be needed for the convergence analysis of each algorithm. Our first two assumptions are required for both algorithms.
Assumption 1.5.
The solution set of problem (1.21), i.e., $\operatorname{SOL}(F,C)$, is nonempty.
Assumption 1.6.
The numerical sequence $\{\beta_{k}\}$ of stepizes is square summable but its sum diverges; that is, it satisfies the following conditions
(1.22) $\sum_{k=0}^{\infty}\beta_{k}$ $=\infty,$
and
(1.23) $\sum_{k=0}^{\infty}\beta_{k}^{2}$ $<\infty.$
The next two assumptions are required in the analysis of Algorithm 1
Assumption 1.7.
The mapping $F$ is paramonotone on $C$ and continuous on $\mathds{R}^{n}$.
Assumption 1.8.
The mapping $F$ is monotone on $C$ and continuous on $\mathds{R}^{n}$.
Assumption 1.9.
Some comments are in order regarding Assumption 1.9. Coerciveness assumptions have been required in previous convergence analyzes of direct methods for variational inequalities, for instance in [34], where a condition stronger than Assumption 1.9 is demanded, and in [17], where Assumption 1.9 is used. Our assumption holds immediately when $F$ is monotone and has a zero $\bar{z}$, in which case we take $\hat{z}=\bar{z}$ and $W=\emptyset$, and also when $F$ is uniformly monotone with $\lim_{t\to\infty}\psi(t)/t=\infty$ (for more details, see [17], where other monotone operators satisfying this assumption are discussed).
We end this section with the following assumption, which is required in the analysis of Algorithm 2.
Assumption 1.10.
The mapping $F$ is monotone on $C$ and continuous on $\mathds{R}^{n}$.
2. Preliminaries
@@
Assume that $K\subset\mathds{R}^{n}$ is a closed convex set. Then, for any $x\in\mathds{R}^{n}$ and any $y\in K$ we have the following:
  1. (i)
    $\langle x-P_{K}(x),y-P_{K}(x)\rangle\leq 0$,
  2. (ii)
@@
If $H\coloneqq\{y\in\mathds{R}^{n}:\left\langle a,y\right\rangle\leq\alpha\}$, with $a\in\mathds{R}^{n}$, $\alpha\in R$, then the orthogonal projection onto $H$ is given by
$P_{H}(x)=x-\frac{\max\{0,\alpha-\left\langle a,x\right\rangle\}}{\|a\|^{2}}a.$
@@
Suppose that $C\subset\mathds{R}^{n}$ is a closed convex set and $U\subset\mathds{R}^{n}$ is an affine subspace. Let $S\subset\mathds{R}^{n}$ be a separating operator for $C$ which satisfies (i)(iii)(i)-(iii) in Definition 1.4. Define $H(x)\subset\mathds{R}^{n}$ as:
(2.1) $H(x)=\begin{cases}K,&{\rm if}\,\,\,x\in C,\\ \{z\in\mathds{R}^{n}:\langle z-P_{S}(x),x-P_{S}(x)\rangle\leq 0\},&{\rm otherwise}.\end{cases}$
@@
Next, we present an explicit instance of the approximate circumcenter operator $T_{\mathbf{S},\mathds{D}}$, defined in (1.18). First, we define a separating operator $S_{i}(y)\subset\mathds{R}^{n}$ for each set $C_{i}$, assumed to be of the form:
(2.3) $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$
@@
(2.4) $S_{i}(y)=\begin{cases}C_{i}&{\rm if}\,\,y\in C_{i},\\ \{z\in\mathds{R}^{n}:\left\langle u_{i},z-y\right\rangle+g_{i}(x)\leq 0\},&{\rm otherwise},\end{cases}$
where $u_{i}$ is a subgradient of $g_{i}$ at $y$. Given ${\mathbf{y}}=(y^{1},\dots,y^{m})\in\mathds{R}^{nm}$, we define ${\mathbf{S}}({\mathbf{y}})\subset\mathds{R}^{nm}$ as
(2.5) ${\mathbf{S}}({\mathbf{y}})=S_{1}(y^{1})\times\dots\times S_{m}(y^{m}).$
We have the following result related to these separating operators.
@@
Suppose that $C_{i}\subset\mathds{R}^{n}$ and $S_{i}(y)\subset\mathds{R}^{n}$ are defined as in (2.3) and (2.4). Then $C_{i}\subset S_{i}(y)$ for any $y\in\mathds{R}^{n}$ and $S_{i}(y)$ is a separating operator for $C_{i}$ at the point $y$ that satisfies conditions (i)-(iii) in Definition 1.4. Also, ${\mathbf{S}}({\mathbf{y}})$ is a separating operator for $C_{1}\times\dots\times C_{m}\subset\mathds{R}^{nm}$ at $\mathbf{y}$}.
@@
Take $C_{i}$ as in (2.3), $T_{\mathbf{S}_{k},\mathbf{D}}$ as in (1.18) and $\mathbf{S}_{k}:=S_{1}(x^{k})\times\cdots\times S_{m}(x^{k})$ where $S_{i}(x^{k})$ is defined in (2.4). By Proposition 2.4 and 2.3, we immediately have $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)=P_{\mathbf{H}_{k}\cap\mathbf{D}}(x)$, where $\mathbf{H}_{k}:=H_{1}(x^{k})\times\cdots\times H_{m}(x^{k})$ and $H_{i}(x^{k})$ is defined in a similar way as $H(x)$ in (2.1) w.r.t. $C_{i}$.
@@
(2.6) $\left\lVert T_{K,U}(x)-y\right\rVert\leq\left\lVert T_{{\rm MAP}}(x)-y\right\rVert,$
for all $x\in U$, for all $y\in K\cap U$, here $T_{K,U}$ is the circumcenter operator as defined in (1.8) and $T_{{\rm MAP}}=P_{U}\circ P_{K}$ is the alternating projection operator with respect to $K$ and $U$.
@@
If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is continuous then $\operatorname{SOL}(T,C)$ is closed.
@@
This follows from the definition of $\operatorname{VIP}(F,C)$ and the continuity of $F$. ∎
@@
Let $F:\mathds{R}^{n}\to\mathds{R}^{n}$ be a continuous and paramonotone operator. Let $\{z^{k}\}$ be a bounded sequence such that all cluster points of $\{z^{k}\}$ belong to $C$. For each $x\in\operatorname{SOL}(F,C)$ define $\gamma_{k}\coloneqq\langle F(x^{k}),z^{k}-x\rangle$. If for some $\bar{x}\in\operatorname{SOL}(F,C)$ there exists a subsequence $\{\gamma_{k_{j}}\}$ of $\{\gamma_{k}\}$ such that $\lim_{j\to\infty}\gamma_{k_{j}}\leq 0$, then there exists a cluster point of $\{z^{k_{j}}\}$ belonging to $\operatorname{SOL}(F,C)$.
@@
Take $\{\varepsilon_{k}\}$, $\{v_{k}\}$ $\subset$ $\mathds{R}_{+}$ and $\mu\in[0,1)$. If the inequalities
(2.7) $\varepsilon_{k+1}\leq\mu\varepsilon_{k}+v_{k},~~k\in\mathds{N},$
hold and $\lim_{k\rightarrow\infty}v_{k}=0$, then $\lim_{k\rightarrow\infty}\varepsilon_{k}=0$.
@@
(2.8) $\operatorname*{dist}(z^{k_{j}+1},K)>\operatorname*{dist}(z^{k_{j}},K),$
@@
We say $w\in\mathds{R}^{n}$ is a Slater point for $\operatorname{SOL}(F,C)$ if
(2.10) $g_{i}(w)<0,$
@@
Let $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ be monotone and continuous and $C$ a closed and convex set. Then $\operatorname{SOL}(F,C)$ is closed and convex.
@@
(2.12) $\operatorname{SOL}(F,C)=\{x\in C:\langle F(y),y-x\rangle\geq 0,~\forall y\in C\}.$
@@
If $\{x^{k}\}$ is a quasi-Fejér convergent to $S$ then:
  1. (i)
@@
  1. (i)
    the sequence $\{\operatorname*{dist}(x^{k},S)\}$ is convergent;
  2. (ii)
@@
Let $g_{i}:\mathds{R}^{n}\rightarrow\mathds{R}$ is convex function for all $i\in\{1,\ldots,m\}$ and $C_{i}\coloneqq\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Define $C\coloneqq\cap_{i=1}^{m}C_{i}$. Assume that there exists a Slater point $w\in C$ such that $g(w)<0$ where $g$ is defined in (4.3). Then, for all $y$ such that $g(y)>0$, we have
(2.14) $\operatorname*{dist}(y,C)\leq\frac{\|y-w\|}{g(y)-g(w)}g(y).$
@@
Take $w_{\lambda}\coloneqq\lambda w+(1-\lambda)y$ with $\lambda=\frac{g(y)}{g(y)-g(w)}$. Note that $\lambda\in(0,1)$. Then
(2.15) $g(w_{\lambda})=g(\lambda w+(1-\lambda)y)\leq\lambda g(w)+(1-\lambda)g(y)=g(y)-\lambda(g(y)-g(w))=0,$
@@
(2.16) $\operatorname*{dist}(y,C)\leq\|y-w_{\lambda}\|=\|y-(\lambda w+(1-\lambda)y)\|=\lambda\|y-w\|=\frac{g(y)}{g(y)-g(w)}\|y-w\|.$
3. A circumcenter method for the variational inequality problem with paramonotone operator
In this section, we consider a circumcenter method, based on the GT method, for solving $\operatorname{VIP}(F,C)$ with a paramonotone operator $F$ and a feasible set of the form $C=\cap_{i=1}^{m}C_{i}$.
Our proposed method is constructed upon Algorithms 1 and 2 proposed by Bello-Cruz and Iusem [17], which are versions of the GT method for point-to-set paramonotone operators. Our approach is to employ a circumcenter-scheme simliarly to what was introduced [15] in order to accelerate the ones in [17].
3.1. The algorithm
We present in Algorithm 1 the one-step circumcenter method for solving $\operatorname{VIP}(F,C)$, with paramonotone $F$ and feasible set $C$ of the form $C=\cap_{i=1}^{m}C_{i}$.
Algorithm 1 Approximate Circumcenter method for $\operatorname{VIP}(F,C)$ with paramonotone $F$
Sequence $\{\beta_{k}\}$ as defined in Assumption 1.6..
Initialization: Choose $x^{0}\in\mathds{R}^{n}$.
for $k=0,1,2,\dots$ do $\triangleright$ Iterative step
  if $F(x^{k})=0$ then
   stop
  end if
  Define $\eta_{k}\coloneqq\max\{1,\left\lVert F(x^{k})\right\rVert\}$.
  Choose a separating operator $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ for the set $C_{i}$.
  Define $S^{i}_{k}\coloneqq S^{i}(x^{k})$ and ${\mathbf{S}}_{k}\coloneqq S^{1}_{k}\times\dots\times S^{m}$.
  Compute the next iterate:
(3.1) $x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right),$
where $T_{\textbf{S}_{k},\textbf{D}}$ is the approximate circumcentered-reflection operator (defined in (1.20)) evaluated at $x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})$, and the $\textbf{S}_{k}$ is determined by $x^{k}$.
end for